OpenStudy (ray1998):

Help please? Will give fan and medal. Assume that each of the n trials is independent and that p is the probability of success on a given trial. Use the binomial probability formula to find P(x). n = 6 x = 3 p = 0.2 p(x) = Thank you in advance :)

OpenStudy (bobo-i-bo):

The binomial distribution formula: if there are n independent trials and the probability of success is p, then: \[P(x)=\left(\begin{matrix}n \\ x\end{matrix}\right)p^x(1-p)^{1-x}\]

OpenStudy (ray1998):

Soo, p(3) = (6/3)p^3)(1 - p)^1-3 Something like that? @Bobo-i-bo

OpenStudy (bobo-i-bo):

yup ^_^

OpenStudy (ray1998):

Okay, thank you!!!!

OpenStudy (bobo-i-bo):

np! :)

OpenStudy (reemii):

(6/3) is not 6÷3, but \(C^6_3 = 20\).

OpenStudy (ray1998):

Ahh, okay! How would I type that though?

OpenStudy (reemii):

(6/3) doesn't look like anything but 6÷3. You can just use (6C3) for example. Do you know the formula for this quantity?

OpenStudy (ray1998):

Alrighty, thank you :) Um, not really, no.

OpenStudy (bobo-i-bo):

http://www.mathwords.com/c/combination_formula.htm

OpenStudy (reemii):

^ yup

OpenStudy (ray1998):

Thank you guys :)

OpenStudy (ray1998):

I applied the formula. (6C3) = 6*5*4*3/3*2*1*0 = 0 Am I right?